Identify if the following statements are predicate logic. Give a domain of discourse for each propositional function.

(x+2)^2 is a prime number?


1
Expert's answer
2022-04-06T09:26:33-0400

'(x+2)2 is a prime number' is a predicate logic and the domain of discourse here is the set of numbers {"\sqrt{2}-2" ,"\sqrt{3}-2" , "\sqrt{5}-2," "\sqrt{7}-2" ...}.